David needs R30 835, 42 to buy a second hand trailer.how lobg will it take him to save towsrds this amount , if he deposit R25 000 now into a savings account earning 10% interest per year compounded weekly


1
Expert's answer
2021-10-06T13:15:58-0400

Compound interest


"A=P(1+\dfrac{r}{n})^{nt}"

Solve for "t"


"(1+\dfrac{r}{n})^{nt}=\dfrac{A}{P}"

"nt=\dfrac{\ln(\dfrac{A}{P})}{\ln(1+\dfrac{r}{n})}"

"t=\dfrac{\ln(\dfrac{A}{P})}{n\ln(1+\dfrac{r}{n})}"

Given "A=30 835.42, P=25000, r=0.1, n=52"


"t=\dfrac{\ln(\dfrac{30 835.42}{25000})}{52\ln(1+\dfrac{0.1}{52})}\approx2.0999 (years)"

"t\approx109.2\ weeks"

It will take him 110 weeks to save towards this amount.